OpenStudy (anonymous):

3(x + 1) = −2(x − 1) + 6 ????

OpenStudy (thurscon):

Expanded form: 3x+3=8-2x x=1

OpenStudy (lanre):

3(x+1) = -2(x-1)+6 3x + 3 = - 2x + 8 3x + 2x = 8 - 3 5x = 5 x = 1